Adjacent sides of a rectangle differ by 5 ft. If the area of the rectangle is 84 sqft., what is the length of the smaller side? Thanks!!
Answer by stanbon(75887)

Adjacent sides of a rectangle differ by 5 ft. If the area of the rectangle is 84 sqft., what is the length of the smaller side?
------------------
Let smaller side be "x":
Then larger side is "x+5":
----
Area = (1/2)x(x+5)
84 = (1/2)(x^2+5x)
x^2 + 5x - 168 = 0
---
x = [-5 +- sqrt(25-4*-168)]/2
---
x = [-5 +-sqrt(697)]/2
---
Positive solution:
---------------------------
x = [21.4]/2
---
x = 10.7 ft
x+5 = 15.7 ft.
==================
cheers,
Stan H.
==========